# A Fractal-based Complex Belief Entropy for Uncertainty Measure in Complex Evidence Theory

### v1

@@ Methods
Shannon entropy Pal et al.’s entropy [37] Deng entropy [35] SU measurement
@@
Proof IV.1.
When there exists a elemnt in the FoD satisfying $\mathbb{M}({{e}_{i}})=1$, the minimum of FCB entropy is $\mathbb{E}_{FCB}^{\min}(\mathbb{M})=0$. In Definition III.3, the maximum of FCB entropy is defined as $\mathbb{E}_{FCB}^{\max}(\mathbb{M})=\log({{2}^{\left|\Theta\right|}}-1)=\log({{2}^{n}}-1)$. Therefore, the range of FCB entropy is $[0,\log({{2}^{n}}-1)]$. So FCB entropy doesn’t satisfy Property IV.3.
Property IV.4 (Monotonicity).
$\small GBe{{l}_{1}}({{P}_{i}})\geq GBe{{l}_{2}}({{P}_{i}}),$ (46)
it must satisfy the following relationship:
$\small{{\mathbb{E}}_{FCB}}({{\mathbb{M}}_{1}})\leq{{\mathbb{E}}_{FCB}}({{\mathbb{M}}_{2}}).$ (47)
Proof IV.2.
Negation is a good way to analyze problems in evidence theory. Considering events from the negative side, some problems that cannot be concluded from the positive side can be dealt with. As the number of negative iterations increases, the amount of information in evidence theory decreases and the uncertainty increases. In this paper, the negative idea is used to prove the monotonicity of FCB entropy.
Case 1. When CBBA is defined on the complex plane. Compared with the traditional entropy model, FCB entropy is used to measure the uncertainty of CBBA, and can deal with information in complex form. The monotonicity is verified by negative thinking. In CET, a new CBBA exponential negation method was proposed [55]. It provides a method to verify the monotonicity of FCB entropy. Consider a FoD $\Theta=\{{{e}_{1}},{{e}_{2}}\}$, select the initial value that is easy to calculate. And after 10 negative iterations, the results of CBBA in CET changing with the number of negations are shown in Table III. The calculation results of FCB entropy are shown in Table IV. It can be seen from Table IV that FCB entropy increases with the number of negation. So when FCB entropy is used on CBBA, it satisfies monotonicity.
TABLE III: CBBA changing with the number of negative iterations.
CBBA
Times $\mathbb{M}(e_{1})$ $\mathbb{M}(e_{2})$ $\mathbb{M}(\Theta)$
0 0.200+0.300i 0.500-0.100i 0.300-0.200i
1 0.211+0.050i 0.289-0.050i 0.500
2 0.240+0.012i 0.260-0.012i 0.500
3 0.248+0.003i 0.252-0.003i 0.500
4 0.249+0.001i 0.251-0.001i 0.500
5 0.250+0.001i 0.250-0.001i 0.500
6 0.250+0.000i 0.250-0.000i 0.500
7 0.250+0.000i 0.250-0.000i 0.500
8 0.250+0.000i 0.250-0.000i 0.500
9 0.250+0.000i 0.250-0.000i 0.500
10 0.250+0.000i 0.250-0.000i 0.500
TABLE IV: Results of each iteration of FCB entropy.
@@
TABLE V: BBA changing with the number of negative iterations.
Times
BBA 0 1 2 3 4 5 6 7 8 9 10
$m(e_{1})$ 0.3333 0.1667 0.0833 0.0417 0.0208 0.0104 0.0052 0.0026 0.0013 0.0001 0.0000
$m({e}_{2})$ 0.3333 0.1667 0.0833 0.0417 0.0208 0.0104 0.0052 0.0026 0.0013 0.0001 0.0000
$m(\Theta)$ 0.3333 0.6667 0.8333 0.9167 0.9583 0.9781 0.9896 0.9948 0.9974 0.9998 1.0000
Case 2. According to Axiom III.1, as a generalization of FB entropy, when CBBA degenerates to BBA, FCB entropy degenerates into FB entropy. Luo proposed a negative method of BBA. The negative direction is the direction of neglect. Similar to the process in Case 1. Consider a FoD $\Theta=\{{{e}_{1}},{{e}_{2}}\}$, select the initial value that is easy to calculate. And after 10 negative iterations, the results of BBA in DSET changing with the number of negations are shown in the Table V. The calculation results of entropies are shown in Table VI. In the meantime, the change trend of FCB entropy and other previous entropies in evidence theroy with the number of iterations are shown in Figure 6. According to Figure 6, it’s clear to see that FB entropy [60] and FCB entropy increase with the number of iterations increasing, while Deng entropy [35], Pal et al.s’ entropy [37], Zhou et al.’s entropy [64] and Cui et al.s’ entropy [62] first increase and then decrease with the number of iterations increasing. That is, FB entropy and FCB entropy meet the requirements of monotonicity, while Deng entropy, Pal et al.s’ entropy, Zhou et al.’s entropy and Cui et al.s’ entropy does not meet the requirements of monotonicity.
@@
TABLE VI: Results of each iteration of different entropies.
Times
Measures 0 1 2 3 4 5 6 7 8 9 10
Deng entropy [35] 2.1133 2.3083 2.1375 1.9500 1.8105 1.7189 1.6624 1.6289 1.6095 1.5986 1.5924
FB entropy [60] 1.3921 1.5420 1.5746 1.5824 1.5843 1.5848 1.5849 1.5850 1.5850 1.5850 1.5850
Pal et al.s’ entropy [37] 1.9183 1.9182 1.6500 1.4138 1.2500 1.1461 1.0835 1.0470 1.0261 1.0144 1.0078
Zhou et al.’s entropy [64] 1.8728 1.8274 1.5364 1.2888 1.1192 1.0126 0.9486 0.9113 0.8900 0.8782 0.8715
Cui et al.s’ entropy [62] 1.4721 1.5068 1.2558 1.0283 0.8687 0.7671 0.7056 0.6696 0.6490 0.6374 0.6309
FCB entropy 1.3921 1.5420 1.5746 1.5824 1.5843 1.5848 1.5849 1.5850 1.5850 1.5850 1.5850
@@
$\small{{\mathbb{E}}_{FCB}}({{\mathbb{M}}_{\Psi}})={{\mathbb{E}}_{FCB}}({{\mathbb{M}}_{\Gamma}})+{{\mathbb{E}}_{FCB}}({{\mathbb{M}}_{\Upsilon}}).$ (48)
@@
Proof IV.3.
Given $\psi\in{{2}^{\left|\Psi\right|}}$, $\tau\in{{2}^{\left|\Gamma\right|}}$ and $\gamma\in{{2}^{\left|\Upsilon\right|}}$, and the relationship of the three FoDs is $\Psi=\Gamma\times\Upsilon$. According to Definition III.1, the following equation is derived
@@
$\small\left|\mathbb{M}_{F}^{\Psi}(\psi)\right|=\left|\mathbb{M}_{F}^{\Gamma}(\tau)\right|\times\left|\mathbb{M}_{F}^{\Upsilon}(\gamma)\right|,$ (51)
@@
$\small\mathbb{C}o{{m}_{F}}(\psi)=\frac{\left|\mathbb{M}_{F}^{\Gamma}(\tau)\right|\times\left|\mathbb{M}_{F}^{\Upsilon}(\gamma)\right|}{\sum\limits_{{{\tau}_{i}}\in{{2}^{\Gamma}};{{\gamma}_{i}}\in{{2}^{\Upsilon}}}{\left|\mathbb{M}_{F}^{\Gamma}({{\tau}_{i}})\right|\times\left|\mathbb{M}_{F}^{\Upsilon}({{\gamma}_{i}})\right|}}.\$ (52)
It is proved that the joint CBBA and FCBBA are continuous in the framework. In addition, it is also known that Shannon entropy is additive, and FCB entropy is similar to Shannon entropy. It is easy to prove that FCB entropy is additive. An example is given below to verify that FCB entropy satisfies additivity.
Example IV.1.
Given two independent FoD $X=\{{{x}_{1}},{{x}_{2}}\}$ and $Y=\{{{y}_{1}},{{y}_{2}}\}$, two CBBAs are defined on the two FoDs, respectively. Let $Z=X\times Y$ be a joint FoD. And the complex mass functions are
${{\mathbb{M}}^{X}}:$ ${{\mathbb{M}}^{X}}({{x}_{1}})=0.2+0.1i,{{\mathbb{M}}^{X}}({{x}_{2}})=0.5+0.1i,$
${{\mathbb{M}}^{X}}({{x}_{1}},{{x}_{2}})=0.3-0.2i,$
${{\mathbb{M}}^{Y}}:$ ${{\mathbb{M}}^{Y}}({{y}_{1}})=0.3+0.2i,{{\mathbb{M}}^{Y}}({{y}_{2}})=0.2+0.1i,$
${{\mathbb{M}}^{Y}}({{y}_{1}},{{y}_{2}})=0.5-0.3i.$
Using the normalized $\mathbb{C}o{{m}_{F}}$ to calculate the FCB entropy, it can be obtained that
${{\mathbb{E}}_{FCB}}({{\mathbb{M}}_{Z}})=2.8317={{\mathbb{E}}_{FCB}}({{\mathbb{M}}_{X}})+{{\mathbb{E}}_{FCB}}({{\mathbb{M}}_{Y}}),$
which indicates that FCB entropy satisfies additivity.
Property IV.6 (Subadditivity).
@@
$\small{{\mathbb{E}}_{FCB}}({{\mathbb{M}}_{\Psi}})\leq{{\mathbb{E}}_{FCB}}({{\mathbb{M}}_{\Gamma}})+{{\mathbb{E}}_{FCB}}({{\mathbb{M}}_{\Upsilon}}).$ (53)
Proof IV.4.
Whether the two FoDs $\Gamma$ and $\Upsilon$ are independent will have an impact on the entropy of the combined FoD $\Psi$, which is mainly divided into the following two cases. Case 1. If CBBAs in $\Gamma$ and $\Upsilon$ are independent of each other, it can be known from Property 5 that ${{\mathbb{E}}_{FCB}}({{\mathbb{M}}_{\Psi}})={{\mathbb{E}}_{FCB}}({{\mathbb{M}}_{\Gamma}})+{{\mathbb{E}}_{FCB}}({{\mathbb{M}}_{\Upsilon}})$.
@@
Property IV.7 (Time complexity).
Let a CBBA $\mathbb{M}$ be defined on a FoD $\Theta=\{{{e}_{1}},{{e}_{2}},\cdots,{{e}_{n}}\}$ which contains $n$ elements. The measurement of FCB entropy can be divided into four steps. The total time complexity of each step and the algorithm is summarized in Table VII.
TABLE VII: Time complexity of each step in the propoed FCB entropy
[b]
Steps Time complexity Description
Step 1 $O(n{{2}^{n}})$ Calculate FCBBA based on CBBA
Step 2 $O({{2}^{n}})$ Calculate the modulus of FCBBA.
Step 3 $O({{2}^{n}})$ Calculate $\mathbb{C}o{{m}_{F}}$ of each focal element.
Step 4 $O({{2}^{n}})$ Finally, calculate FCB entropy.
Step 1-4 $O(n{{2}^{n}})$ Time complexity.
TABLE VII: Time complexity of each step in the propoed FCB entropy
According to Table VII, it can be seen that the computational complexity of FCB entropy is similar to that of FB entropy [60] and JS entropy [39], and the computational complexity of some methods is higher than that of FCB entropy. Therefore, the computational complexity of FCB entropy is still within an acceptable range.
Property IV.8 (Discord and non-specificity).
@@
$\small\mathbb{E}_{FCB}^{\mathcal{D}}(\mathbb{M})=-\sum\limits_{i=1}^{n}{\mathbb{C}omCBet({{e}_{i}})\log\left(\mathbb{C}omCBet({{e}_{i}})\right)},$ (54)
@@
$\small\mathbb{C}omCBet({{e}_{i}})=\frac{\left|CBet({{e}_{i}})\right|}{\sum\limits_{i=1}^{n}{\left|CBet({{e}_{i}})\right|}},$ (55)
@@
Among the existing methods, there are entropies that can be directly decomposed into discord and non-specificity from the formula, such as Deng entropy and JS entropy. However, similar to FB entropy, the measurement of FCB entropy on discord and non-specificity is not directly obtained from the formula. FCB entropy originates from the production process of CPBT. When CBBA is converted to probability through CPBT, there is only uncertainty caused by discord in the framework. Therefore, the definitions of discord and non-specificity in FCB entropy are as above.
In Definition IV.1, $\mathbb{E}_{FCB}^{\mathcal{D}}(\mathbb{M})$ represents the uncertainty caused by singletons, excluding the uncertainty caused by set allocation in evidence theory. So it is reasonable to represent the discord in FCB entropy. Total uncertainty minus the uncertainty caused by discord is the uncertainty caused by non-specificity. Therefore, FCB entropy satisfies Property IV.8.
An example is given below to show the relationship between the discord and non-specificity of FCB entropy.
@@
Then according to Definition III.1, the FCBBA can be obtained
@@
V Experiments
The generation of FCBBA includes the process of assigning multi-element focal elements to its subsets, reflecting the impact of the intersection of different events on the system uncertainty measurement. Therefore, FCB entropy can be used to reflect the uncertainty in CBBA caused by the intersection of different events. Next, several specific examples are used to illustrate.
Example V.1.
The purpose of the question is to find the target element. Suppose there are two different sources of evidence, that is, two CBBAs are defined on a same FoD $X=\{{{x}_{1}},{{x}_{2}},{{x}_{3}},{{x}_{4}}\}$, the results of the two CBBAs are given as below.
${{\mathbb{M}}_{1}}:{{\mathbb{M}}_{1}}(\{{{x}_{1}},{{x}_{2}}\})=0.4+0.2i,{{\mathbb{M}}_{1}}(\{{{x}_{3}},{{x}_{4}}\})=0.6-0.2i,$
${{\mathbb{M}}_{2}}:{{\mathbb{M}}_{2}}(\{{{x}_{1}},{{x}_{3}}\})=0.4+0.2i,{{\mathbb{M}}_{1}}(\{{{x}_{2}},{{x}_{3}}\})=0.6-0.2i.$
It can be seen from the two CBBAs that the four elements in the first evidence source may be the target elements, and the corresponding elements in the two complex mass functions given are different from each other. In the second evidence source, there are only three elements as possible target elements, and there is a same element ${{x}_{3}}$ in the two complex mass functions given. Therefore, it is intuitively judged that the conflict degree of ${{\mathbb{M}}_{1}}$ is higher than that of ${{\mathbb{M}}_{2}}$, that is, the uncertainty of ${{\mathbb{M}}_{2}}$ should be less than that of ${{\mathbb{M}}_{1}}$.
In order to compare the effect of FCB entropy and the previously proposed entropies on multi-element intersection, the previous entropies are extended to CBBA using Definition LABEL:def2.16. The calculation results of each entropy are shown in Table VIII.
TABLE VIII: Results of different entropies in Example V.1.
Entropy
$\mathbb{M}$ Pal et al. [37] Deng [35] Zhou et al. [64] Cui et al. [62] FCB
$\mathbb{M}_{1}$ 1.9787 2.5636 2.2029 2.5636 2.5636
$\mathbb{M}_{2}$ 1.9787 2.5636 2.2029 2.0827 2.2529
From Table VIII, it is obvious that Pal et al.’s entropy, Deng entropy, Zhou et al. s’entropy did not take into account the change in the amount of information caused by the intersection of focal elements, so the entropy of the two evidence sources is the same, resulting in information loss. Only the uncertainty of FCB entropy measurement meets the requirement that ${{\mathbb{M}}_{1}}$ is greater than ${{\mathbb{M}}_{2}}$. Logically, ${{\mathbb{M}}_{1}}$ should contain more information because the uncertainty is greater than ${{\mathbb{M}}_{2}}$. Therefore, the newly proposed FCB entropy is more effective in reducing information loss.
Here, another example is given to verify its effectiveness in measuring uncertain information.
Example V.2.

### v2

@@ Methods
Shannon entropy [67] Pal et al.’s entropy [49] Deng entropy [51] SU measurement
@@
TABLE III: Results of each iteration of FCB entropy.
@@
Proof 3. When there exists a elemnt in the FoD satisfying $\mathbb{M}({{e}_{i}})=1$, the minimum of FCB entropy is $\mathbb{E}_{FCB}^{\min}(\mathbb{M})=0$. In Definition 21, the maximum of FCB entropy is defined as $\mathbb{E}_{FCB}^{\max}(\mathbb{M})=\log({{2}^{\left|\Theta\right|}}-1)=\log({{2}^{n}}-1)$. Therefore, the range of FCB entropy is $[0,\log({{2}^{n}}-1)]$. So FCB entropy doesn’t satisfy Property 3.
Property 4 (Monotonicity).
$GBe{{l}_{1}}({{P}_{i}})\geq GBe{{l}_{2}}({{P}_{i}}),$ (43)
it must satisfy the following relationship:
${{\mathbb{E}}_{FCB}}({{\mathbb{M}}_{1}})\leq{{\mathbb{E}}_{FCB}}({{\mathbb{M}}_{2}}).$ (44)
Proof 4. Negation plays a pivotal role in evidence theory, offering a valuable perspective for problem analysis. By examining events from a negative standpoint, it becomes possible to address issues that may not be resolvable from a positive perspective. As the count of negative iterations increases, evidence theory posits that the amount of information diminishes while uncertainty escalates. This paper employs the concept of negation to demonstrate the monotonicity of FCB entropy.
Case 1. When CBBA is defined over the complex plane, FCB entropy is leveraged to gauge the uncertainty of CBBA, effectively handling complex-form information. The monotonicity is established through negative reasoning. In the context of CET, a novel exponential negation method for CBBA has been introduced [39], offering a technique to affirm the monotonicity of FCB entropy. Consider a Frame of Discernment (FoD) $\Theta=\{{{e}_{1}},{{e}_{2}}\}$ and select an initial value conducive to calculation. The corresponding computations of FCB entropy are presented in Table III. Observations from Table III indicate that FCB entropy ascends with the increasing number of negations, thereby confirming the monotonicity of FCB entropy when applied to CBBA. However, it can be seen from Figure 4 that QB entropy does not satisfy monotonicity.
Case 2. According to Axiom 1, as a generalization of FB entropy, when CBBA degenerates to BBA, FCB entropy degenerates into FB entropy. Luo proposed a negative method of BBA. The negative direction is the direction of neglect. Similar to the process in Case 1. Consider a FoD $\Theta=\{{{e}_{1}},{{e}_{2}}\}$, select the initial value that is easy to calculate. The change trend of FCB entropy and other previous entropies in evidence theroy with the number of iterations are shown in Figure 4. According to Figure 4, it’s clear to see that FB entropy [61] and FCB entropy increase with the number of iterations increasing, while Deng entropy [51], Pal et al.’s entropy [49], Zhou et al.’s entropy [69] and Cui et al.’s entropy [70] first increase and then decrease with the number of iterations increasing. That is, FB entropy and FCB entropy meet the requirements of monotonicity, while Deng entropy, Pal et al.’s entropy, Zhou et al.’s entropy and Cui et al.’s entropy does not meet the requirements of monotonicity.
@@
${{\mathbb{E}}_{FCB}}({{\mathbb{M}}_{\Psi}})={{\mathbb{E}}_{FCB}}({{\mathbb{M}}_{\Gamma}})+{{\mathbb{E}}_{FCB}}({{\mathbb{M}}_{\Upsilon}}).$ (45)
@@
Proof 5. Given $\psi\in{{2}^{\left|\Psi\right|}}$, $\tau\in{{2}^{\left|\Gamma\right|}}$ and $\gamma\in{{2}^{\left|\Upsilon\right|}}$, and the relationship of the three FoDs is $\Psi=\Gamma\times\Upsilon$. According to Definition 19, the following equation is derived
@@
$\left|\mathbb{M}_{F}^{\Psi}(\psi)\right|=\left|\mathbb{M}_{F}^{\Gamma}(\tau)\right|\times\left|\mathbb{M}_{F}^{\Upsilon}(\gamma)\right|,$ (48)
@@
$\mathbb{C}o{{m}_{F}}(\psi)=\frac{\left|\mathbb{M}_{F}^{\Gamma}(\tau)\right|\times\left|\mathbb{M}_{F}^{\Upsilon}(\gamma)\right|}{\sum\limits_{{{\tau}_{i}}\in{{2}^{\Gamma}};{{\gamma}_{i}}\in{{2}^{\Upsilon}}}{\left|\mathbb{M}_{F}^{\Gamma}({{\tau}_{i}})\right|\times\left|\mathbb{M}_{F}^{\Upsilon}({{\gamma}_{i}})\right|}}.\$ (49)
It is proved that the joint CBBA and FCBBA are continuous in the framework. In addition, it is also known that Shannon entropy is additive, and FCB entropy is similar to Shannon entropy. It is easy to prove that FCB entropy is additive. An example 4 is given in appendix to verify that FCB entropy satisfies additivity.
Property 6 (Subadditivity).
@@
${{\mathbb{E}}_{FCB}}({{\mathbb{M}}_{\Psi}})\leq{{\mathbb{E}}_{FCB}}({{\mathbb{M}}_{\Gamma}})+{{\mathbb{E}}_{FCB}}({{\mathbb{M}}_{\Upsilon}}).$ (50)
Proof 6. Whether the two FoDs $\Gamma$ and $\Upsilon$ are independent will have an impact on the entropy of the combined FoD $\Psi$, which is mainly divided into the following two cases.
Case 1. If CBBAs in $\Gamma$ and $\Upsilon$ are independent of each other, it can be known from Property 5 that ${{\mathbb{E}}_{FCB}}({{\mathbb{M}}_{\Psi}})={{\mathbb{E}}_{FCB}}({{\mathbb{M}}_{\Gamma}})+{{\mathbb{E}}_{FCB}}({{\mathbb{M}}_{\Upsilon}})$.
@@
Property 7 (Time complexity).
Let a CBBA $\mathbb{M}$ be defined on a FoD $\Theta=\{{{e}_{1}},{{e}_{2}},\cdots,{{e}_{n}}\}$ which contains $n$ elements. The measurement of FCB entropy can be divided into four steps. The total time complexity of each step and the algorithm is summarized in Table IV.
[b] Steps Time complexity Description Step 1 $O(n{{2}^{n}})$ Calculate FCBBA based on CBBA Step 2 $O({{2}^{n}})$ Calculate the modulus of FCBBA. Step 3 $O({{2}^{n}})$ Calculate $\mathbb{C}o{{m}_{F}}$ of each focal element. Step 4 $O({{2}^{n}})$ Finally, calculate FCB entropy. Steps 1-4 $O(n{{2}^{n}})$ Time complexity.
TABLE IV: Time complexity of each step in the propoed FCB entropy
  • Explanation of notations in the table:
  • 1
    $n$: The number of baic event in the frame of discernment $\Theta$.
  • 2
    $O(f\left(n\right))$: The maximum time complexity.
According to Table IV, it can be seen that the computational complexity of FCB entropy is similar to that of FB entropy [61] and JS entropy [52], and the computational complexity of some methods is higher than that of FCB entropy. Therefore, the computational complexity of FCB entropy is still within an acceptable range.
Property 8 (Discord and non-specificity).
@@
$\mathbb{E}_{FCB}^{\mathcal{D}}(\mathbb{M})=-\sum\limits_{i=1}^{n}{\mathbb{C}omCBet({{e}_{i}})\log\left(\mathbb{C}omCBet({{e}_{i}})\right)},$ (51)
@@
$\mathbb{C}omCBet({{e}_{i}})=\frac{\left|CBet({{e}_{i}})\right|}{\sum\limits_{i=1}^{n}{\left|CBet({{e}_{i}})\right|}},$ (52)
@@
Among the existing entropy measures, some can explicitly decompose discord and non-specificity directly from their formulas, such as Deng entropy and JS entropy. However, akin to FB entropy, FCB entropy does not straightforwardly yield measures of discord and non-specificity from its formula. The FCB entropy’s genesis lies in the Complex Pignistic Belief Transformation (CPBT) process. When CBBA is translated into probability via CPBT, the resulting uncertainty is solely attributed to discord within the framework. Hence, the definitions of discord and non-specificity in FCB entropy are as previously stated.
In Definition 22, $\mathbb{E}_{FCB}^{\mathcal{D}}(\mathbb{M})$ signifies the uncertainty due to singletons, excluding the uncertainty stemming from set allocation in evidence theory. This makes it a fitting representation for discord in FCB entropy. The total uncertainty minus the discord-induced uncertainty leaves the uncertainty due to non-specificity. Consequently, FCB entropy adheres to Property 8.
An example is provided to illustrate the interplay between discord and non-specificity within FCB entropy.
@@
Then according to Definition 19, the FCBBA can be obtained
@@
V APPLICATION
In this section, we initially employ illustrative example to highlight the benefits of FCB entropy. Subsequently, we apply the pattern recognition algorithm from [60] and the information fusion processing algorithm from [43], substituting the entropy measure with FCB entropy for experimental validation. The outcomes underscore the effectiveness of the proposed FCB entropy.
Figure 6: Line chart of calculation results of different entropy in Example 3.
V-A Discuss from numerical examples
The construction of FCBBA entails the allocation of multi-element focal elements to subsets, capturing the effect of event intersections on the system’s uncertainty assessment. Consequently, FCB entropy serves as an indicator of the uncertainty within CBBA that arises from overlapping events. To elucidate this concept, a toy example will be presented.
Example 3.